OpenStudy (anonymous):

Find the equation of the line that passes through the points (-2, 3) and (1, -6). Write your answer in Standard Form. Answer -3x – y = 3 -3x – y = -3 3x + y = 3 3x + y = -3

OpenStudy (anonymous):

Always go back to y = mx+b

OpenStudy (anonymous):

you have to find the slope first by using y2-y1/x2-x1

OpenStudy (anonymous):

@Bailey.Nae do you know how to find your slope given these two points?

OpenStudy (anonymous):

once you find the slope put it into y=mx+b and change it to standard form

OpenStudy (anonymous):

slope in -3

OpenStudy (anonymous):

The best way here is to use the options

OpenStudy (anonymous):

3x + y = -3

OpenStudy (anonymous):

Also, ax+by=c is the standard from (in general)

OpenStudy (anonymous):

Yep. Okay you've got m=-3 and two points you can choose from. Either (-2,3) or (1,-6) It DOESN'T MATTER which point you decide to work with. I'll pick (1,-6) to work with. And then there's m=-3 Your first step is to solve for b. y=mx+b Plug the values in and solve. Once you have your b-value, plug your m-value and your b-value in to y=mx+b and you'll have your equation for your line.